If you have a 3% solution and add equil amounts of a 6% solution, what is the % solution?
THANK YOU!

Found 2 solutions by tommyt3rd, Alan3354:
Answer by tommyt3rd(5050)

x(0.03)+x(0.06)=x(0.09)

new solution is 9%


:)

Answer by Alan3354(69443)

If you have a 3% solution and add equil amounts of a 6% solution, what is the % solution?
------------
It's the average, (3 + 6)/2 = 4.5%